OpenStudy (anonymous):

express all radian measures in terms of pi express -5pi/6 radians in degrees?

OpenStudy (anonymous):

\[= -5\pi/6\] Alright, in order to convert it into degrees, you need to multiply that to 180/pi. So, \[(\frac{ -5\pi }{ 1 }) \times (\frac{ 180 }{ \pi }) \] Pi crosses out. Then you're left with:\[-5\times180= -900 degrees\] Hope this helps!

OpenStudy (anonymous):

Oh shoots, I forgot the divided by 6, hahaha just divide the whole thing by 6, it'll then come out to -150 degrees :)

OpenStudy (anonymous):

thank you so much but one question it has to be in terms of pi would it be -150 pi?

OpenStudy (anonymous):

Hmmmmm, lemme think on that.

OpenStudy (anonymous):

oh you crossed out the pi

OpenStudy (anonymous):

but it needs to be in terms of pi i been trying to do th

OpenStudy (anonymous):

Wouldn't it already be in Pi? since you're starting with 5pi/6?

OpenStudy (anonymous):

That's the way I learnt it, I may be wrong though :P

OpenStudy (anonymous):

the directions are so confusing thats what it says originally the question says express all radian measures for problems 1-4 in terms of pi #2 express -5pi/6 radians in degrees im so confused

OpenStudy (anonymous):

lol i dont doubt the way you learned it the way you did is very clear :)

OpenStudy (anonymous):

No this one is right. The question IS in terms of pi, they must be asking for the degrees.

OpenStudy (anonymous):

Because 150 degrees in terms of pi is 5pi/6, I checked on my circle chart thing. :P

OpenStudy (anonymous):

So I have no idea what the other answer could be.

OpenStudy (anonymous):

lol im not good in math at all but you must be right because somebody looked at this problem as well they couldnt do it another way either lol

OpenStudy (anonymous):

u are right